a certain number is 2 more than 3 times another number. if their sum is 14, find the numbers.
Answer by macston(5194)

.
x=a certain number; y=another number
.
x+y=14
x=14-y
.
x=3y+2
14-y=3y+2
12=4y
3=y ANSWER 1: One of the numbers is 3.
.
x=14-y=14-3=11 ANSWER 2: The other number is 11.